calculate the number of molecules of C8H8 containing 1.44 * 10^25 C atoms
1
Expert's answer
2011-10-07T08:24:50-0400
N - number of C8H8 molecules:

N (C8H8) = (1.44 * 10^25)/8 = 0.18 * 10^25
= 18 * 10^23 molecules